The instaneous value of voltage in an a. c circuit at any time t seconds is given by v=100sin(50πt-0.523) V. Find



1:The peak to peak voltage, the frequency and the phase angle



2:the voltage when t=0



3:the voltage when t=8ms



4:the times in the first cycle when the voltage is 60V and -40V



5:the first time when the voltage is maximum

1
Expert's answer
2022-02-07T15:16:15-0500

Part 1

peak to peak voltage "=100\u00d72=200V"

"2\pi\>f=50\pi\implies\>f=25Hz"


Phase angle"=0.523" radians ; lagging


Part 2


"V=100\>Sin(-0.523)"

"=100\u00d7-0.49948"

"=-49.948V"


Part 3

"V=100" Sin "(50\pi×8×10^{-3}-0.523)"


"=100\u00d70.6697"

"=66.97V"


Part 4

"60=100" Sin "(50\pi\>t-0.523)"

"50\pi\>t-0.523=" Sin"^{-1}\frac{60}{100}"

"50\pi\>t-0.523=0.6435"

"t=7.426ms"


"-40=100" Sin "(50\pi\>t-0.523)"

"50\pi\>t-0.523=" Sin"^{-1}(\frac{-40}{100})"

"50\pi\>t-0.523=-0.4115"

"t=0.7097ms"


Part 5

"100=100" Sin"(50\pi\>t-0.523)"

"50\pi\>t-0.523=" Sin"^{-1}1"

"50\pi\>t-0.523=1.5708"

"t=13.33ms"